### v1

@@ 6 Discussion
In this study, we investigated the spike-adding mechanisms in a three-dimensional neuronal model with fast, slow, and super-slow timescales, inspired by the FitzHugh-Nagumo (FHN) model driven by low-frequency periodic inputs. The FHN model, a two-dimensional simplification of the Hodgkin-Huxley model, effectively captures neuronal excitability and the characteristic large excursions known as spikes. In addition to being mathematically tractable, the FHN model introduces parabolic bursts, a class of bursting that has received less theoretical attention compared to others.
@@
Our analysis, based on both numerical simulations and geometric singular perturbation theory, highlighted how the number of spikes within each burst can be regulated by tuning the frequency and amplitude of the external input. We found and characterized key bifurcations and mechanisms such as the folded saddle and folded node canards, that drive spike addition and shape the boundaries between regions with different spike counts. We identified six regions representing different behaviors in the FHN model with periodic input, based on the input’s frequency $\omega$ and amplitude $E$. A summary of the results is presented in Figure 3. The analytically calculated boundaries between these regions correspond to saddle-node bifurcations of folded equilibria on $\mathcal{L}_{-}$ (or $\mathcal{L}_{+}$) - the left (right) knee of the critical manifold $\mathcal{M}$ or correspond to the changes in stability from node to saddle on $\mathcal{L}_{-}$ (or $\mathcal{L}_{+}$). We did not analyze the system’s dynamics inside Region I (where the amplitude $E$ is small) because there was no bursting there. We focused instead on Regions II and III (Figure 3; $E$ of moderate values) and discussed two key spike-adding mechanisms.
@@
One way to add spikes in the burst is through a folded saddle canard; this situation occurred in both regions II and III. Another mechanism, which only appears in region II, is to create spikes through a folded node canard. Additionally, we showed that spike initiation at the folded node was influenced by both the canard and the delayed-Hopf mechanisms [25, 27], and that the relative influence of these mechanisms shifted as the input frequency changed. Furthermore, we demonstrated how these spike-adding mechanisms organize the boundaries in parameter space that separate regions with different spike counts (Figure 2 and Figure 3– right panel) by numerically calculating the $L^{2}$ norms of trajectories and estimating the delay-to-spike within each burst.
Together, these findings provide deeper insights into the dynamics of parabolic bursting, a less-studied class of bursting behavior, and offer a framework for understanding similar phenomena in other excitable systems. Future research will focus on exploring the dynamics in regions IV, V, and VI of the parameter space in Figure 3, where the amplitude $E$ is high and the system exhibits “held-small amplitude oscillations” as defined by [18]. We also plan to extend in the future this analysis to other models of neuronal excitability.
@@ 7 Appendix
7.1 Numerical simulations for bifurcation diagram
The bifurcation diagrams for spike counts (Figures 12) and $L^{2}$ norms (Figure 9) were obtained with a parameter sweep of the forcing amplitude, $E$, and frequency, $\omega$. Each simulation was initialized at the equilibrium point of the unforced problem ($E=0$), $(x,y)\approx(-1.1994,-1.4993)$. The simulation was then burned in for two input periods to allow transient solutions to drop out.
@@ 7.1 Numerical simulations for bifurcation diagram
7.2 Series approximations to invariant manifolds of the saddle
We compute series expansions for the stable and unstable manifolds of the folded saddle. The stable eigenvector of the saddle (see Proposition 3.8) is given by
@@ 7.2 Series approximations to invariant manifolds of the saddle
In the $\theta\times u$ plane, this corresponds to a slope of $-\lambda_{S\ell,1}/(2\delta)$, which is finite for $\delta\neq 0$. So, we write the stable manifold of the saddle, $W^{s}_{S\ell}$, as a graph with a series expansion of $u$ in terms of $\theta$. But first, we translate the saddle to the origin with $\hat{\theta}=\theta-\theta_{S\ell}$, so that we can write
(27) $u=a_{1}\hat{\theta}+a_{2}\hat{\theta}^{2}+a_{3}\hat{\theta}^{3}+a_{4}\hat{\theta}^{4}+a_{5}\hat{\theta}^{5}+\mathcal{O}(\hat{\theta}^{6}),$
@@
(28) $\dfrac{du}{d\hat{\theta}}=a_{1}+2a_{2}\hat{\theta}+3a_{3}\hat{\theta}^{2}+4a_{4}\hat{\theta}^{3}+5a_{5}\hat{\theta}^{4}+\mathcal{O}(\hat{\theta}^{5}).$
@@
(29) $\dfrac{du}{d\hat{\theta}}=\frac{\mu\cos\hat{\theta}-\sqrt{R_{\delta}^{2}-\mu^{2}}\sin\hat{\theta}-G(u)}{\delta u(u-2)}.$
@@
(30) $\dfrac{du}{d\hat{\theta}}=b_{0}+b_{1}\hat{\theta}+b_{2}\hat{\theta}^{2}+b_{3}\hat{\theta}^{3}+b_{4}\hat{\theta}^{4}+\mathcal{O}(\hat{\theta}^{5}),$
@@
(31) $a_{1}$ $=b_{0}\left(a_{1},\dots,a_{5};\delta,R_{\delta},\mu\right)$
@@
To promote a curve tangent to the stable eigenvector, we use the initial guess
@@
The same procedure is used to calculate the local unstable manifold of the folded saddle, $W^{u}_{S\ell}$, except the initial guess provided to ‘fsolve()’ promotes a curve tangent to the unstable eigenvector (see Proposition 3.8),
@@
7.3 Saddle invariant manifold coefficients
@@
Choosing the subtraction term gives exactly $a_{1}=-\lambda_{S\ell,1}/(2\delta)$ as expected for the stable manifold. The other solution gives $a_{1}=-\lambda_{S\ell,2}/(2\delta)$ as expected for the unstable manifold (see Proposition 3.8).

### v2

@@ 7 Discussion
In this study, we investigated the spike-adding mechanisms in a three-dimensional neuronal model with fast, slow, and super-slow timescales, inspired by the FitzHugh-Nagumo (FHN) model driven by low-frequency periodic inputs, and tested on the Morris-Lecar (ML) as well. The FHN and ML models, two-dimensional reductions of the Hodgkin-Huxley model, effectively capture neuronal excitability and the characteristic large excursions known as spikes.
@@
Our analysis, based on both numerical simulations and geometric singular perturbation theory, highlighted how the number of spikes within each burst can be regulated by tuning the frequency and amplitude of the external input. We found and characterized key bifurcations and mechanisms such as the folded saddle and folded node canards, that drive spike addition and shape the boundaries between regions with different spike counts. We identified six regions representing different behaviors in the FHN model with periodic input, based on the input’s frequency $\omega$ and amplitude $E$. A summary of the results is presented in Figure 3. The analytically calculated boundaries between these regions correspond to saddle-node bifurcations of folded equilibria on $\mathcal{L}_{-}$ (or $\mathcal{L}_{+}$), the left (right) knee of the critical manifold $\mathcal{M}$, or correspond to the changes in stability from folded node to folded saddle on $\mathcal{L}_{-}$ (or $\mathcal{L}_{+}$). We did not analyze the system’s dynamics inside Region I (where the amplitude $E$ is small) because there was no bursting there. We focused instead on Regions II and III (Figure 3; $E$ of moderate values) and discussed two key spike-adding mechanisms.
@@
One way to add spikes in the burst is through a folded saddle canard; this situation occurred in both regions II and III. Another mechanism, which only appears in region II, is to create spikes through a folded node canard. Furthermore, we demonstrated how these spike-adding mechanisms organize the boundaries in parameter space that separate regions with different spike counts (Figure 2 and Figure 3– right panel) by numerically calculating the $L^{2}$ norms of trajectories and estimating the delay-to-spike within each burst.
We speculate that pitchfork and transcritical bifurcations of maximal canards ([32, Figure 8.12]), arising at specific ratios of the folded node eigenvalues, may govern the escape dynamics of solutions as they pass through the folded node. To explore this, we computed eigenvalue ratios at the folded node across the parameter space and overlaid their level sets on the spike-adding regions (figures are not shown). We observed partial alignment between low-integer ratio contours and cusp locations in the spike-adding diagrams, particularly for the first few transitions. Although these patterns became less clear at larger (biologically realistic) time-scale parameters, they support the hypothesis that canard bifurcations could influence the underlying structure of spike-adding dynamics, especially in the extreme (outside the biological relevant) multiple-timescale regime. A more systematic study of this relationship remains a promising direction for future work.
The bursting patterns analyzed in our paper do not fit into any of the three main known bursting types – square, elliptic or parabolic. In particular, these bursts do not satisfy the mathematical definition of the parabolic bursting since the fast subsystem of the fast/slow/super-slow system is only one-dimensional and cannot generate spikes on its own. However, we observed in our numerical experiments (not shown here) that the bursting patterns of the forced FHN model exhibit a parabolic-burster–like property. Shortly, their inter-spike intervals increase towards the end of the bursting phase, akin to those reported in classical parabolic bursting [3]. These bursts also show an elliptic-burst-like feature. Indeed, both the FHN and the ML models with constant input exhibit bistability (admittedly, in a very small parameter range), with the stable limit cycle surrounding the stable equilibrium point e.g., see Figure 12,C). When this fast(er) dynamics is coupled with the slow dynamics of the driving force, the trajectory slowly drifts between the spiking and quiet phases like in the elliptic bursting. Thus, while the bursts investigated in our work are neither parabolic nor elliptic, they still share some attributes with each of them.
Our numerical results provide compelling evidence of a consistent structure in the spike-adding transitions and enhance the current understanding of spike-adding mechanisms. Future research will focus on exploring the dynamics in regions IV, V, and VI of the parameter space in Figure 3, where the amplitude $E$ is high and the system exhibits “held-small amplitude oscillations” as defined by [52]. We acknowledge that a full characterization of the bifurcation structure of spike-adding dynamics is an important theoretical question that merits further investigation. This is especially true in systems with three distinct timescales and canard-induced transitions, where standard bifurcation tools may be insufficient. Developing and applying appropriate techniques to address this challenge is a natural next step in our research. We also plan to extend in the future this analysis to other models of neuronal excitability.
Acknowledgment
This work was funded by NSF-RTG award 1840260 (PM), The Stanley-UI Foundation Support Organization (RC), and NSF award 2037828 and The Simons Foundation MPS–TSM–00008005 (ZA).
@@ 8 Appendix
8.1 Numerical simulations for spike-adding diagram
The bifurcation diagrams for spike counts (Figures 12) and $L^{2}$ norms (Figure 9) were obtained with a parameter sweep of the forcing amplitude, $E$, and frequency, $\omega$. Each simulation was initialized at the equilibrium point of the unforced problem ($E=0$), $(x,y)\approx(-1.1994,-1.4993)$. We first ran the simulation for two input periods that were then excluded from the analysis, to allow transient solutions to drop out.
@@ 8.1 Numerical simulations for spike-adding diagram
  • For the Morris-Lecar simulations of Figure 12: $(f,E)\in[0.5,10]\times[10,70]$ with step-sizes of $5\times 10^{-3}$ and $7.5\times 10^{-2}$ for $f$ and $E$, respectively.
8.2 Series approximations to invariant manifolds of the folded saddle
We compute series expansions for the stable and unstable manifolds of the folded saddle. The stable eigenvector of the folded saddle (see Proposition 3) is given by
@@ 8.2 Series approximations to invariant manifolds of the folded saddle
In the $\theta\times u$ plane, this corresponds to a slope of $-\lambda_{S\ell,1}/(2\delta)$, which is finite for $\delta\neq 0$. So, we write the stable manifold of the folded saddle, $W^{s}_{S\ell}$, as a graph with a series expansion of $u$ in terms of $\theta$. But first, we translate the folded saddle to the origin with $\hat{\theta}=\theta-\theta_{S\ell}$, so that we can write
$u=a_{1}\hat{\theta}+a_{2}\hat{\theta}^{2}+a_{3}\hat{\theta}^{3}+a_{4}\hat{\theta}^{4}+a_{5}\hat{\theta}^{5}+\mathcal{O}(\hat{\theta}^{6}),$ (29)
@@
$\dfrac{du}{d\hat{\theta}}=a_{1}+2a_{2}\hat{\theta}+3a_{3}\hat{\theta}^{2}+4a_{4}\hat{\theta}^{3}+5a_{5}\hat{\theta}^{4}+\mathcal{O}(\hat{\theta}^{5}).$ (30)
@@
$\dfrac{du}{d\hat{\theta}}=\frac{\mu\cos\hat{\theta}-\sqrt{R_{\delta}^{2}-\mu^{2}}\sin\hat{\theta}-G(u)}{\delta u(u-2)}.$ (31)
@@
$\dfrac{du}{d\hat{\theta}}=b_{0}+b_{1}\hat{\theta}+b_{2}\hat{\theta}^{2}+b_{3}\hat{\theta}^{3}+b_{4}\hat{\theta}^{4}+\mathcal{O}(\hat{\theta}^{5}),$ (32)
@@
$a_{1}$ $=b_{0}\left(a_{1},\dots,a_{5};\delta,R_{\delta},\mu\right)$ (33)
@@
To generate a curve tangent to the stable eigenvector, we use the initial guess
@@
The same procedure is used to calculate the local unstable manifold of the folded saddle, $W^{u}_{S\ell}$, except the initial guess provided to ‘fsolve()’ promotes a curve tangent to the unstable eigenvector (see Proposition 3),
@@
8.3 Folded saddle invariant manifold coefficients
@@
Choosing the subtraction term gives exactly $a_{1}=-\lambda_{S\ell,1}/(2\delta)$ as expected for the stable manifold. The other solution gives $a_{1}=-\lambda_{S\ell,2}/(2\delta)$ as expected for the unstable manifold (see Proposition 3).
@@
8.4 Morris-Lecar equations and parameters
The system of differential Equations in Eq. (27) is supplemented by the steady-state equations,
$m_{\infty}(V)$ $=\frac{1}{2}\left[1+\tanh((V-V_{1})/V_{2})\right]$
${\tau_{n}}(V)$ $=1/\cosh((V-V_{3})/(2V_{4})),$
and voltage-dependent time constant for potassium activation
${n_{\infty}}(V)=\frac{1}{2}\left[1+\tanh((V-V_{3})/V4)\right].$
The parameters were chosen to induce subcritical Hopf bifurcations and are given in Table 1.
Table 1: Parameters used (see Eq. (27)) for Morris-Lecar simulations with constant and periodic input.
Parameter $\phi$ $g_{\text{Ca}}$ $V_{3}$ $V_{4}$ $E_{\text{Ca}}$ $E_{\text{K}}$ $E_{\text{L}}$ $g_{\text{K}}$ $g_{\text{L}}$ $V_{1}$ $V_{2}$ $C_{M}$
Value 0.075 4.4 2 30 120 -84 -60 8 2 -1.2 18 20